OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. Can somebody show me how to do this?

OpenStudy (anonymous):

since the new line should be parallel to the given this means they have the same slope in this case -3 general eqn of a line is: y=ax+b a=-3 a point (2,6) belongs to the line so just substitute the values in the general form you have x y and a u just need b 6=-3*2+b b=12 so the eqn is y=-3x+12

OpenStudy (anonymous):

Nevermind, figured it out! :))